Assume that good A and good B are related goods and QB=1691-400PB+6PA-6Y. Suppose that PB=0.1 Birr, PA=0.3 Birr and Y (income) =10 Birr. Then compute.
- Price elasticity of demand for good B.
Income elasticity of demand for good B and
- explain nature of the good.
- Cross price elasticity of demand and state the nature of the goods whether they are substitute, complementary or unrelated.
1
Expert's answer
2021-08-10T10:31:17-0400

Solution:

First derive the value of QB:

QB = 1691 – 400(0.1) + 6(0.3) – 6(10) = 1691 – 40 + 1.8 – 60 = 1691 + 1.8 – 40 – 60 = 1592.8

QB = 1592.8

a.). Price elasticity of demand for good B:

PEd = "\frac{\triangle Q}{\triangle PB }\times \frac{PB}{Q }"


"PEd = -400\times \frac{0.1}{1592.8 } = -0.025"

The demand for good B is price inelastic, that is demand is not sensitive to price changes.


b.). Income elasticity of demand for good B:

Income elasticity of demand (YEd) = "\frac{\triangle Q}{\triangle I }\times \frac{I}{Q }"


"-6\times \frac{10}{1592.8 } = -0.038"

The demand is income inelastic. Therefore, the good is inferior since it has a negative income elasticity of demand, which means that demand falls as income increases.


c.). Cross price elasticity of demand and state the nature of the goods:

Cross price elasticity of demand of good B with respect to good A:


Cross price elasticity of demand = "\frac{\triangle Q}{\triangle PA }\times \frac{PA}{Q }"


"6\times \frac{0.3}{1592.8 } = 0.0012"

The goods are substitutes since the value is greater than zero.